Hello I'm having difficult time understanding this problem?
4 1/4 times 3 3/8
My answer comes to 14 34/32, after I divided 459 by 32.
but the correct answer is 14 11/32
Where is this remainder of 11 found, how do I find it?
Please explain.
Thanks,
-Garrett Young

Answer by Alan3354(21550)

4 1/4 times 3 3/8
-------------
= 17/4 times 27/8
= 459/32
14*32 = 448
459 - 448 = 11
= 14 11/32